Solve for x:
-7[4x-(x-5)]+2x=3-x
I keep coming up with 16/9 - but the book sez I'm wrong... what am I doing wrong?

Answer by stanbon(48557)

-7[4x-(x-5)]+2x=3-x
-7[4x-x+5]+2x=3-x
-7[3x+5]+2x=3-x
-21x-35+2x=3-x
18x=-38
x=-19/9
Cheers,
Stan H.